OpenStudy (anonymous):

The function f(x) = 14(3)^x represents the growth of a dragonfly population every year in a remote swamp. Rose wants to manipulate the formula to an equivalent form that calculates seven times a year, not just once a year. Which function is correct for Rose's purpose, and what is the new growth rate?

OpenStudy (anonymous):

f(x) = 2(3)^x; growth rate 200% f(x) = 2(1.17)^7x; growth rate 117% f(x) = 2(3)^x, growth rate 2% f(x) = 2(1.17)^x; growth rate 2%

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

@amoodarya @chmvijay @Darrius @elephantine

OpenStudy (anonymous):

@whalexnuker... did you ever figure this out!? i have the same problem!

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@amoodarya @chmvijay @Darrius @elephantine

OpenStudy (anonymous):

@deniseeeex3 I don't remember... I'm not enrolled in that course anymore :/ sorry...